19.) A ball is thrown horizontally from the top of a building 130 m high. The ball strikes the ground 59 m horizontally from the point of release. What is the speed of the ball just before it strikes the ground ?


1
Expert's answer
2021-09-27T14:13:31-0400

The time of falling can be found from the kinematic equation:


"h = \dfrac{gt^2}{2}"

where "h = 130m" if the height of the builing, and "g = 9.81m\/s^2" if the gravitational acceleration.

Thus, obtain:


"t = \sqrt{\dfrac{2h}{g}}"

The speed in veritical direction by the time "t" is (motion with uniform acceleration):


"v_y = v_{y0} + gt = gt = \sqrt{2gh}"

where "v_{y0} = 0m\/s" is the initial vertical speed.

The speed in horizontal direction by the time "t" is (motion with constant speed):


"v_x = d/t = d\sqrt{\dfrac{g}{2h}}"

where "d = 59m" is the range.

The total speed is then:


"v = \sqrt{v_x^2 + v_y^2} = \sqrt{2gh + \dfrac{d^2g}{2h}} = \sqrt{9.81\cdot \left(2\cdot 130 + \dfrac{59^2}{2\cdot 130}\right)} \approx 51.8m/s"

Answer. 51.8 m/s.